# WaveletDiff: Multilevel Wavelet Diffusion For Time Series Generation

### v1

@@ Abstract
Time series are ubiquitous in many applications that involve forecasting, classification and causal inference tasks, such as healthcare, finance, audio signal processing and climate sciences. Still, large, high-quality time series datasets remain scarce. Synthetic generation can address this limitation; however, current models confined either to the time or frequency domains struggle to reproduce the inherently multi-scaled structure of real-world time series. We introduce WaveletDiff, a novel framework that trains diffusion models directly on wavelet coefficients to exploit the inherent multi-resolution structure of time series data. The model combines dedicated transformers for each decomposition level with cross-level attention mechanisms that enable selective information exchange between temporal and frequency scales through adaptive gating. It also incorporates energy preservation constraints for individual levels based on Parseval’s theorem to preserve spectral fidelity throughout the diffusion process. Comprehensive tests across six real-world datasets from energy, finance, and neuroscience domains demonstrate that WaveletDiff consistently outperforms state-of-the-art time-domain and frequency-domain generative methods on both short and long time series across five diverse performance metrics. For example, WaveletDiff achieves discriminative scores and Context-FID scores that are $3\times$ smaller on average than the second-best baseline across all datasets.
@@ 1 Introduction
Time series data arises in diverse practical settings, including healthcare Lee et al. (2019a); van der Schaar Lab (2019), finance Sezer et al. (2020); Özbayoğlu et al. (2020), climate sciences Dinku (2019); Climate Change AI (2021), audio processing Mitra & Zualkernan (2025) and engineering Susto et al. (2020); Lei et al. (2020); Carvalho et al. (2022). Due to various constraints, acquiring sufficiently high-quality labeled time-series datasets remains a challenge Wang et al. (2021); Desai et al. (2025). The problem may be mitigated through synthetic time series generation, which also offers promising solutions for data augmentation Wen et al. (2021); Le Guennec et al. (2018); Ryu et al. (2023), privacy preservation Wang et al. (2020); Jordon et al. (2024); Nosowsky & Giordano (2006), forecasting Taga et al. (2025) and simulations Nikolenko (2021); El Emam et al. (2022).
@@
Current time series generation methods predominantly operate either directly in the time domain or frequency domain, and come with different advantages and limitations. Time-domain approaches, including those based on GANs Yoon et al. (2019); Pei et al. (2021); EskandariNasab et al. (2024), autoregressive Salinas et al. (2020) and diffusion models Lim et al. (2023); Narasimhan et al. (2024); Sikder et al. (2024) are well-suited for modeling local temporal patterns, but struggle with long-term dependencies and preservation of important spectral characteristics. To address time-domain induced limitations, recent approaches have increasingly leveraged frequency-domain analysis, often along with temporal modeling Tian et al. (2020); Chi et al. (2024); Crabbé et al. (2024); Huang et al. (2024). These methods are of relevance since many real-world time series tend to exhibit higher localization in the frequency rather than the time domain. Representative methods include FourierFlow Alaa et al. (2021), which applies normalizing flows to Fourier representations, DiffusionTS Yuan & Qiao (2024), which combines Fourier decompositions with diffusion models, and various frequency-enhanced transformers Zhou et al. (2022a; b); Xu et al. (2024); Yi et al. (2023) that use both spectral and temporal analyses. However, these approaches typically process time and frequency domain information either in a separate manner or impose trade-offs between temporal resolution and spectral coherence. They are also not able to simultaneously capture both local and global time and spectral patterns, which is crucial for synthesizing realistic time series.
@@
Wavelet transforms represent a natural approach to address the above issue by creating a multi-resolution representation that simultaneously captures both temporal and spectral information Mallat (1989); Cohen (2001). Unlike the Fourier transform, which capture global frequency properties, wavelets maintain temporal localization while also providing useful decompositions into multiple frequency bands Rioul & Flandrin (1992); Daubechies (1988). This results in a highly versatile time-frequency hierarchical representation Mallat (1989). As a result, wavelet-based analyses have been used with success for various signal processing applications including speech recognition, financial trends analysis, image processing, and biomedical signal analysis Daubechies (1992); Vetterli & Kovačević (1995); Burrus et al. (1998). Despite these results, a handful of known wavelet-based approaches for time series generation have failed to provide improvements over Fourier-based methods Takahashi & Mizuno (2024); Kazemi & Meidani (2022); Zhao et al. (2018). This may be attributed to the fact that, almost exclusively, the methods treat wavelet coefficients as image structures and then follow-up by applying standard image generation techniques such as convolutional neural networks or image-based diffusion models. While potentially useful for data-poor applications and highly specialized time series, indirect time series $\to$ wavelet $\to$ image conversion methods in general suffer from pattern distortions caused by noninvertible image features.
@@
A more adequate approach based on wavelet decompositions is to run diffusion models directly in the wavelet domain, which is a new direction proposed in this work. For diffusion, methods such as denoising diffusion probabilistic models (DDPMs) Ho et al. (2020) which have demonstrated remarkable success in image Dhariwal & Nichol (2021), audio Kong et al. (2021), and text generation Austin et al. (2021), are considered state-of-the-art for time-series generation. However, these diffusion models are tailor-made for highly specific time-series formats (e.g., audio or financial data), and may not be suitable for other modalities. This motivates us to implement a new wavelet-space diffusion model, termed WaveletDiff, which is universally applicable as it inherently respects different multi-level structures. Unlike frequency-domain approaches, WaveletDiff also captures temporal patterns at different scales simultaneously. Our key innovations lie in running forward diffusion processes for each wavelet level individually and in parallel, following fine-tuned exponential noise mechanisms and using dedicated level-transformer denoising networks combined with a cross-level attention mechanism that enables information exchange between different decomposition levels. This design preserves the hierarchical nature of wavelet representations while allowing the model to learn complex inter-scale dependencies crucial for realistic time series generation.
@@
4. A new evaluation metric based on the Dynamic Time-Warping distance and extensive comparative analysis of both short and long time series datasets, including ETTh1, ETTh2, Stocks, Exchange Rate, fMRI, EEG Zhou et al. (2021); Lai et al. (2018); Roesler (2013). The results show significant performance gains of WaveletDiff compared to Fourier and time-based methods, which are roughly three-fold on average for discriminative and Context-FID scores.
5. The first empirical evaluation of reproducibility of diffusion models for time series, akin to recent efforts reported for images Zhang et al. (2024b); Li et al. (2024); Kadkhodaie et al. (2024).
@@ 2 Related Work
Time Series Generation with Diffusion Models. Early generative AI methods for time series focused on conditional generation tasks such as forecasting and imputation Rasul et al. (2021); Tashiro et al. (2021); Li et al. (2022); Yang et al. (2024), while recent approaches target unconditional time series generation Shen & Kwok (2023); Barancikova et al. (2025). The above methods employ various architectural choices including RNNs, transformers, and specialized denoising networks to handle the sequential nature of temporal data Kong et al. (2021), and almost exclusively operate in the time domain. This limits their ability to capture global and local spectral properties.
@@
Frequency Domain Approaches for Time Series. Recent works have demonstrated that real-world time series are more localized in the frequency domain, making spectral diffusion more effective than time diffusion Crabbé et al. (2024). Various frequency-based approaches include lightweight models using complex-valued operations Xu et al. (2024), frequency-enhanced transformers combining discrete Fourier transform (DFT) with attention mechanisms Zhou et al. (2022b; a), and specialized MLP architectures for frequency learning Yi et al. (2023). Additional methods incorporate spectral filtering Zhang et al. (2024a), multi-resolution frequency analysis Wang et al. (2024), and normalizing flows in Fourier domain Alaa et al. (2021). Additional unconditional generation approaches include interpretable diffusion models that combine trend and seasonality components Yuan & Qiao (2024) and latent diffusion models that operate in compressed latent spaces for more efficient generation Qian et al. (2024). These methods often require separate processing pipelines for temporal and frequency components, limiting their ability to simultaneously capture multi-scale temporal-spectral relationships.
@@
Wavelet-Based Time Series Modeling. Wavelet transforms provide multi-resolution time-frequency representation capabilities Mallat (1989); Daubechies (1992); Addison (2017) and have been extensively used in time series analysis Percival & Walden (2000); Sang (2013); Patrik et al. (2015). In generative modeling, wavelets have shown promise across various domains through direct coefficient processing Phung et al. (2022); Hu et al. (2023); Guth et al. (2022). For time series forecasting, wavelets have been employed to enhance traditional forecasting models through their multi-resolution time-frequency analysis capabilities Zhou et al. (2025); Sasal et al. (2022); Arabi et al. (2024); Schlüter & Deuschle (2010). For generation tasks, existing methods predominantly convert wavelet coefficients to image representations for processing with standard computer vision techniques Takahashi & Mizuno (2024); Kazemi & Meidani (2022). However, this indirect approach may not fully exploit the hierarchical multi-scale structure of wavelet decompositions, where each level captures distinct temporal and spectral characteristics.
@@ 3.2 Wavelet-Space Diffusion Framework
The denoising network $f_{\theta}$ uses dedicated transformers for each wavelet level. We adopt Adaptive Layer Normalization (AdaLN) Peebles & Xie (2022) as the normalization layer. For level $l$, the coefficients $\bm{C}^{(l)}_{t}$ are processed through a specialized transformer,
@@ 4.1 Experimental Settings
We compare WaveletDiff to several state-of-the-art time series generation methods, including FourierDiffusion Crabbé et al. (2024), Diffusion-TS Yuan & Qiao (2024), TimeGAN Yoon et al. (2019), and SigDiffusion Barancikova et al. (2025).
@@
We use six real-world datasets to evaluate our method, covering energy, finance, and neuroscience domains. ETTh1 and ETTh2 Zhou et al. (2021) are electricity transformer datasets containing oil temperature and six power load features recorded hourly from 2016 to 2018. Stocks is a multivariate financial time series dataset containing historical Google stock market data with price and volume features from 2004 to 2019. Exchange Rate Lai et al. (2018) contains daily exchange rates of eight countries from 1990 to 2016. fMRI is the NetSim dataset containing simulated BOLD time series data for evaluating network modeling methods in functional magnetic resonance imaging. EEG Roesler (2013) contains multichannel electroencephalogram recordings that measure brain electrical activity over time, offering information about neural dynamics and cognitive processes. For more details, refer to Appendix B.1.
@@
We evaluate generation quality using five complementary metrics. The discriminative score measures similarity between real and generated samples by training a binary classifier to distinguish them Yoon et al. (2019). The predictive score assesses the utility of synthetic data for forecasting real sequences using mean absolute error. Context-Fréchet inception distance (Context-FID) Paul et al. (2022) quantifies distributional distance using TS2Vec Yue et al. (2022) embeddings following Yuan & Qiao (2024). The correlational score evaluates temporal dependencies by comparing cross-correlation matrices. Additionally, we propose a novel metric, termed the Dynamic Time Warping Jensen-Shannon Distance (DTW-JS distance), which combines the temporal alignment features of DTW with the JS divergence for distributional comparison. DTW aims to capture optimal temporal alignments between two time series sequences $x$ and $y$ by minimizing
@@ 4.2 Short Sequence Time Series Generation
We follow the evaluation setup of TimeGAN Yoon et al. (2019) and Diffusion-TS Yuan & Qiao (2024) to assess generation quality against baseline models. All datasets are segmented into sequences of length $24$ using a sliding window with stride $1$. For evaluation, we generate samples matching the size of the original training data for each dataset to ensure fair evaluation. Training configurations and times, as well as model complexity are discussed in Appendices B.4 and B.5.
@@ 4.4 Ablation Study
To validate the effectiveness of different architectural components of WaveletDiff, we conduct an ablation study comparing our full model against four variants: (1) Predicting coefficients rather than noise: Instead of predicting noise as in standard DDPM, this variant directly predicts the wavelet coefficients themselves, following the approach in Diffusion-TS Yuan & Qiao (2024) which suggests this method outperforms noise prediction. (2) Removing cross-attention: This variant disables information exchange between different wavelet decomposition levels, but maintains level-specific transformer models. (3) Using Cosine instead of Exponential schedules: This variant replaces exponential with cosine noise scheduling during diffusion. (4) Using DDIM sampling: This variant uses deterministic DDIM rather than DDPM sampling during inference. The results in Table 7 in the Appendix reveal that cross-level attention is the universally most critical architectural component.
@@
Inspired by recent diffusion model reproducibility studies Zhang et al. (2024b); Li et al. (2024); Kadkhodaie et al. (2024), we examine whether this phenomenon extends to time series generation. To this end, we train model pairs with architectural variations, generate samples from identical Gaussian noise using deterministic DDIM sampling, and conclude that time series diffusion models exhibit reproducibility across all representation domains (see Appendix D).

### v2

@@ Abstract
Time series are ubiquitous in many applications that involve forecasting, classification and causal inference tasks, such as healthcare, finance, audio signal processing and climate sciences. Still, large, high-quality time series datasets remain scarce. Synthetic generation can address this limitation; however, current models confined either to the time or frequency domains struggle to reproduce the inherently multi-scaled structure of real-world time series. We introduce WaveletDiff, a novel framework that trains diffusion models directly on wavelet coefficients to exploit the inherent multi-resolution structure of time series data. The model combines dedicated transformers for each decomposition level with cross-level attention mechanisms that enable selective information exchange between temporal and frequency scales through adaptive gating. It also incorporates energy preservation constraints for individual levels based on Parseval’s theorem to preserve spectral fidelity throughout the diffusion process. Comprehensive tests across six real-world datasets from energy, finance, and neuroscience domains demonstrate that WaveletDiff consistently outperforms state-of-the-art time-domain and frequency-domain generative methods on both short and long time series across five diverse performance metrics. For example, WaveletDiff achieves discriminative scores and Context-FID scores that are $3\times$ smaller on average than the second-best baseline across all datasets. Our code is available at https://github.com/GarlicWang/WaveletDiff.
@@ 1 Introduction
Time series data arises in diverse practical settings, including healthcare lee2019dynamic; vanderSchaar2019survival, finance sezer2020financial; ozbayoglu2020deep, climate sciences dinku2019challenges; climatechange2021gaps, audio processing mitra2025music and engineering susto2020predictive; lei2020survey; carvalho2022challenges. Due to various constraints, acquiring sufficiently high-quality labeled time-series datasets remains a challenge wang2021deep; desai2025timeseries. The problem may be mitigated through synthetic time series generation, which also offers promising solutions for data augmentation wen2021time; forestier2017generating; ryu2023simpsi, privacy preservation wang2020partgan; jordon2022synthetic; nosowsky2006hipaa, forecasting taga2025timepfn and simulations nikolenko2021synthetic; elEmam2020synthetic.
@@
Current time series generation methods predominantly operate either directly in the time domain or frequency domain, and come with different advantages and limitations. Time-domain approaches, including those based on GANs yoon2019time; pei2021generatingrealworldtimeseries; eskandarinasab2024seriesgantimeseriesgeneration, autoregressive salinas2020deepar and diffusion models lim2023regulartimeseriesgenerationusing; 10.5555/3692070.3693584; sikder2024transfusiongeneratinglonghigh are well-suited for modeling local temporal patterns, but struggle with long-term dependencies and preservation of important spectral characteristics. To address time-domain induced limitations, recent approaches have increasingly leveraged frequency-domain analysis, often along with temporal modeling tian2020tfgantimefrequencydomain; chi2024rfdiffusionradiosignalgeneration; crabbe2024time; huang2024generative. These methods are of relevance since many real-world time series tend to exhibit higher localization in the frequency rather than the time domain. Representative methods include FourierFlow alaa2021generative, which applies normalizing flows to Fourier representations, DiffusionTS yuan2024diffusionts, which combines Fourier decompositions with diffusion models, and various frequency-enhanced transformers zhou2022film; zhou2022fedformer; xu2023fits; yi2023frets that use both spectral and temporal analyses. However, these approaches typically process time and frequency domain information either in a separate manner or impose trade-offs between temporal resolution and spectral coherence. They are also not able to simultaneously capture both local and global time and spectral patterns, which is crucial for synthesizing realistic time series.
@@
Wavelet transforms represent a natural approach to address the above issue by creating a multi-resolution representation that simultaneously captures both temporal and spectral information mallat1989theory; cohen2001uncertainty. Unlike the Fourier transform, which capture global frequency properties, wavelets maintain temporal localization while also providing useful decompositions into multiple frequency bands rioul1992timefrequency; daubechies1988timefrequency. This results in a highly versatile time-frequency hierarchical representation mallat1989theory. As a result, wavelet-based analyses have been used with success for various signal processing applications including speech recognition, financial trends analysis, image processing, and biomedical signal analysis daubechies1992ten; vetterli1995wavelets; burrus1998introduction. Despite these results, a handful of known wavelet-based approaches for time series generation have failed to provide improvements over Fourier-based methods takahashi2024generation; kazemi2022timeseriessynthesismultiscale; 8595010. This may be attributed to the fact that, almost exclusively, the methods treat wavelet coefficients as image structures and then follow-up by applying standard image generation techniques such as convolutional neural networks or image-based diffusion models. While potentially useful for data-poor applications and highly specialized time series, indirect time series $\to$ wavelet $\to$ image conversion methods in general suffer from pattern distortions caused by noninvertible image features.
@@
A more adequate approach based on wavelet decompositions is to run diffusion models directly in the wavelet domain, which is a new direction proposed in this work. For diffusion, methods such as denoising diffusion probabilistic models (DDPMs) ho2020denoising which have demonstrated remarkable success in image dhariwal2021diffusion, audio kong2021diffwave, and text generation austin2021structured, are considered state-of-the-art for time-series generation. However, these diffusion models are tailor-made for highly specific time-series formats (e.g., audio or financial data), and may not be suitable for other modalities. This motivates us to implement a new wavelet-space diffusion model, termed WaveletDiff, which is universally applicable as it inherently respects different multi-level structures. Unlike frequency-domain approaches, WaveletDiff also captures temporal patterns at different scales simultaneously. Our key innovations lie in running forward diffusion processes for each wavelet level individually and in parallel, following fine-tuned exponential noise mechanisms and using dedicated level-transformer denoising networks combined with a cross-level attention mechanism that enables information exchange between different decomposition levels. This design preserves the hierarchical nature of wavelet representations while allowing the model to learn complex inter-scale dependencies crucial for realistic time series generation.
@@
4. A new evaluation metric based on the Dynamic Time-Warping distance and extensive comparative analysis of both short and long time series datasets, including ETTh1, ETTh2, Stocks, Exchange Rate, fMRI, EEG haoyietal-informer-2021; lai2018modelinglongshorttermtemporal; eeg_eye_state_264. The results show significant performance gains of WaveletDiff compared to Fourier and time-based methods, which are roughly three-fold on average for discriminative and Context-FID scores.
5. The first empirical evaluation of reproducibility of diffusion models for time series, akin to recent efforts reported for images pmlr-v235-zhang24cn; li2024understanding; kadkhodaie2024generalization.
@@ 2 Related Work
Time Series Generation with Diffusion Models. Early generative AI methods for time series focused on conditional generation tasks such as forecasting and imputation rasul2021autoregressive; tashiro2021csdi; li2022csdi; NEURIPS2024_5e364212, while recent approaches target unconditional time series generation shen2023timediff; barancikova2025sigdiffusions. The above methods employ various architectural choices including RNNs, transformers, and specialized denoising networks to handle the sequential nature of temporal data kong2021diffwave, and almost exclusively operate in the time domain. This limits their ability to capture global and local spectral properties.
@@
Frequency Domain Approaches for Time Series. Recent works have demonstrated that real-world time series are more localized in the frequency domain, making spectral diffusion more effective than time diffusion crabbe2024time. Various frequency-based approaches include lightweight models using complex-valued operations xu2023fits, frequency-enhanced transformers combining discrete Fourier transform (DFT) with attention mechanisms zhou2022fedformer; zhou2022film, and specialized MLP architectures for frequency learning yi2023frets. Additional methods incorporate spectral filtering zhang2024atfnet, multi-resolution frequency analysis wang2024spectral, and normalizing flows in Fourier domain alaa2021generative. Additional unconditional generation approaches include interpretable diffusion models that combine trend and seasonality components yuan2024diffusionts and latent diffusion models that operate in compressed latent spaces for more efficient generation qian2024timeldmlatentdiffusionmodel. These methods often require separate processing pipelines for temporal and frequency components, limiting their ability to simultaneously capture multi-scale temporal-spectral relationships.
@@
Wavelet-Based Time Series Modeling. Wavelet transforms provide multi-resolution time-frequency representation capabilities mallat1989theory; daubechies1992ten; addison2017wavelet and have been extensively used in time series analysis percival2000wavelet; sang2013review; waveletadvantages2015. In generative modeling, wavelets have shown promise across various domains through direct coefficient processing phung2022wavelet; hu2023neuralwaveletdomaindiffusion3d; NEURIPS2022_03474669. For time series forecasting, wavelets have been employed to enhance traditional forecasting models through their multi-resolution time-frequency analysis capabilities zhou2025multiorderwaveletderivativetransform; Sasal_2022; arabi2024wavemaskmixexploringwaveletbasedaugmentations; Schlter2010UsingWF. For generation tasks, existing methods predominantly convert wavelet coefficients to image representations for processing with standard computer vision techniques takahashi2024generation; kazemi2022timeseriessynthesismultiscale. However, this indirect approach may not fully exploit the hierarchical multi-scale structure of wavelet decompositions, where each level captures distinct temporal and spectral characteristics.
@@ 3.2 Wavelet-Space Diffusion Framework
The denoising network $f_{\theta}$ uses dedicated transformers for each wavelet level. We adopt Adaptive Layer Normalization (AdaLN) Peebles2022DiT as the normalization layer. For level $l$, the coefficients $\bm{C}^{(l)}_{t}$ are processed through a specialized transformer,
@@ 4.1 Experimental Settings
We compare WaveletDiff to several state-of-the-art time series generation methods, including FourierDiffusion crabbe2024time, Diffusion-TS yuan2024diffusionts, TimeGAN yoon2019time, and SigDiffusion barancikova2025sigdiffusions.
@@
We use six real-world datasets to evaluate our method, covering energy, finance, and neuroscience domains. ETTh1 and ETTh2 haoyietal-informer-2021 are electricity transformer datasets containing oil temperature and six power load features recorded hourly from 2016 to 2018. Stocks is a multivariate financial time series dataset containing historical Google stock market data with price and volume features from 2004 to 2019. Exchange Rate lai2018modelinglongshorttermtemporal contains daily exchange rates of eight countries from 1990 to 2016. fMRI is the NetSim dataset containing simulated BOLD time series data for evaluating network modeling methods in functional magnetic resonance imaging. EEG eeg_eye_state_264 contains multichannel electroencephalogram recordings that measure brain electrical activity over time, offering information about neural dynamics and cognitive processes. For more details, refer to Appendix B.1.
@@
We evaluate generation quality using five complementary metrics. The discriminative score measures similarity between real and generated samples by training a binary classifier to distinguish them yoon2019time. The predictive score assesses the utility of synthetic data for forecasting real sequences using mean absolute error. Context-Fréchet inception distance (Context-FID) paul2022psaganprogressiveselfattention quantifies distributional distance using TS2Vec yue2022ts2vecuniversalrepresentationtime embeddings following yuan2024diffusionts. The correlational score evaluates temporal dependencies by comparing cross-correlation matrices. Additionally, we propose a novel metric, termed the Dynamic Time Warping Jensen-Shannon Distance (DTW-JS distance), which combines the temporal alignment features of DTW with the JS divergence for distributional comparison. DTW aims to capture optimal temporal alignments between two time series sequences $x$ and $y$ by minimizing
@@ 4.2 Short Sequence Time Series Generation
We follow the evaluation setup of TimeGAN yoon2019time and Diffusion-TS yuan2024diffusionts to assess generation quality against baseline models. All datasets are segmented into sequences of length $24$ using a sliding window with stride $1$. For evaluation, we generate samples matching the size of the original training data for each dataset to ensure fair evaluation. Training configurations and times, as well as model complexity are discussed in Appendices B.4 and B.5.
@@ 4.4 Ablation Study
To validate the effectiveness of different architectural components of WaveletDiff, we conduct an ablation study comparing our full model against four variants: (1) Predicting coefficients rather than noise: Instead of predicting noise as in standard DDPM, this variant directly predicts the wavelet coefficients themselves, following the approach in Diffusion-TS yuan2024diffusionts which suggests this method outperforms noise prediction. (2) Removing cross-attention: This variant disables information exchange between different wavelet decomposition levels, but maintains level-specific transformer models. (3) Using Cosine instead of Exponential schedules: This variant replaces exponential with cosine noise scheduling during diffusion. (4) Using DDIM sampling: This variant uses deterministic DDIM rather than DDPM sampling during inference. The results in Table 7 in the Appendix reveal that cross-level attention is the universally most critical architectural component.
@@
Inspired by recent diffusion model reproducibility studies pmlr-v235-zhang24cn; li2024understanding; kadkhodaie2024generalization, we examine whether this phenomenon extends to time series generation. To this end, we train model pairs with architectural variations, generate samples from identical Gaussian noise using deterministic DDIM sampling, and conclude that time series diffusion models exhibit reproducibility across all representation domains (see Appendix D).